OpenStudy (anonymous):

You are selling tickets for a high school play. Student tickets cost four dollars and general admission tickets cost six dollars. You sell 525 tickets and collect 2876 dollars. How many of each type of ticket did you sell. Solve using the substitution method( linear systems )

OpenStudy (anonymous):

x + y = 525 and 4x + 6y = 2876. Substitute and solve for each variable.

OpenStudy (campbell_st):

call the tickets s for student and g for general the total tickets sold is s + g = 525 the money collected is 4s + 6g = 2876 so you have 2 equations in 2 unknowns that need to be solved simultaneously s + g = 525 4s + 6g = 2876 hope this helps